OpenStudy (anonymous):

square root of 224 divided by square root of 7 in simplest form

OpenStudy (anonymous):

go to mathway.com it will help

OpenStudy (anonymous):

\[\frac{ \sqrt{224} }{ \sqrt{7} }=\frac{ \sqrt{2^4*2*7} }{ \sqrt{7} }=\frac{ 2^2*\sqrt{2}*\sqrt{7} }{ \sqrt{7}}=?\]

OpenStudy (anonymous):

\[4\sqrt{2}??\]

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

thank you